Find an equation of the line that passes through the given points and has the given slope. Express your answer in slope intercept form. (8,0),m=-3
Answer by scott8148(6628)

using point-slope ___ y - 0 = -3(x - 8)

solve for y to get slope-intercept form (y = mx + b)